# Automatically Identifying Local and Global Circuits with Linear Computation Graphs

### v1

@@ 4 Attributing Intermediate SAE Features
An exciting application of Sparse Autoencoders is that they serve as unsupervised feature extractors in the vast hidden activation space. This opens up opportunities for understanding intermediate activations and local circuit discovery i.e. identifying a subgraph activating a given SAE feature instead of end-to-end circuits.
@@ 4.1 How Transformers Implement In-Bracket Features
We start from a series of In-Bracket features in attention blocks of early layer, which activate on tokens inside of brackets e.g. deactivated [activated] deactivated. These features will demonstrate higher activation in deeper nesting of brackets, imitating the behavior of finite state automata (Bricken et al., 2023) with states of bracket nesting hierarchy. We find an In-Square-Bracket feature and an In-Round-Bracket feature in SAEs trained on layer 1 attention block output, which we call L1A throughout this paper. Since they are at rather early layers, we leverage our Direct Contribution analysis to see how earlier features produce them.
Open-bracket features activate in-bracket ones. Figure 4(a) illustrates a simple two-layer bracket circuit in the wild. We inspect contributions to the In-Square-Bracket feature in a template e.g. "0 0 [1 1 1 [2] 3] 4", at token "1"s, "2", " 3" and " 4". Experiments show that the activation is mainly promoted by an L0M feature activated by token " [". It takes on 104.1%, 102.6% and 314.2% of the In-Square-Bracket feature’s activation respectively at token "1", "2", and " 3", respectively. An average of 83.8% of these contribution comes through attention head 1 of L1A i.e. L1A.H1.
@@ 4.2 Revisiting Induction Behavior from the SAE Lens
Induction Heads (Olsson et al., 2022) are an important type of compositional circuit with two attention layers which try to repeat any 2-gram that occured before i.e. [A][B] … [A] -> [B]. These circuits are believed to account for a majority of in-context learning functionality in large transformers. Compared to massive existing literature in understanding the induction mechanism in the granularity of attention heads (Olsson et al., 2022; Hendel et al., 2023; Ren et al., 2024), inter alia, we seek to present a finer-grained level interpretation of such behavior.
Induction features form a huge feature family. These features are found to be identified by the logit of tokens they enhance through the logit lens (nostalgebraist, 2020). We first study a Capital Induction feature contributing to logits of single capital letters, on a curated input "Video in WebM support: Your browser doesn’t support HTML5 video in WebM." (Figure 5(a)). This feature is activated on the second " Web", and amplifies the prediction of "M", copying its previous occurrence.
@@
Figure 5: "Web"(L0M.1270 and L1M.23399) and "Web" Preceding features (L2A.14876 and L2A.17608) jointly lead to QK attention of an induction head. The "M" feature is copied to the last token for next token prediction.
Upstream Contribution through OV Circuit We notice that a series of "M" features in the residual stream of the first "M" constitute most of the Capital Induction feature’s activation through OV circuits. L0M.88 takes the lead, which contributes 35.0% of the feature activation. Auxiliary features from L0A, L1M and L3M either directly indicate current token as "M", or indicate current token as a single capital letter. Top 7 of the auxiliary features account for another 33.0% of the feature activation. Most of these contributions come from L5A.H1, which we along with a concurrent research (Krzyzanowski2024attention_saes) identify as an induction head.
Upstream Contribution to QK Attention To study how this induction head attend to the first "M", we attribute the attention score to upstream feature pairs. The commonality of top contributors is a " Web" feature attending to a " Web" Preceding feature (i.e. its previous token is " Web"), as shown in Figure 5(b).
Attributing Preceding features We further study how " Web" Preceding features indicate previous tokens. These contributions mainly come through L2A.H2, which we think to be a previous token head. The relatively high attention score for previous token can be attributed to a group of L0A features collecting information from positional embeddings.
@@ 5 Revisiting Indirect Object Identification Circuits from the SAE Lens
Figure 6: In $s_{\text{John}}$, the consecutive entity feature (denoted as A in Figure 6(a)) serves as the key vector for Name Mover Heads to attend to and copy the answer entity to the last token’s residual stream. Such mechanism does not work in $s_{\text{Mary}}$ because the correct answer is no longer a consecutive entity (i.e. the entity present after the token and). See Appendix E for a detailed interpretation for these two examples.
For end-to-end circuits in GPT2-small, we choose to investigate a task called Indirect Object Identification (IOI) (Wang et al., 2023) with Hierarchical Attribution. For instance, GPT2 is able to predict " Mary" following the prompt "When Mary and John went to the store, John gave the bag to". We call this prompt $s_{\text{Mary}}$ since it starts with " Mary" and a variant $s_{\text{John}}$ with a swap in the first two names i.e. "When John and Mary went to the store, John gave the bag to". The answer to both prompts is " Mary", which GPT2 is able to predict. Existing literature studying this problem does not distinguish between these two templates. Through the lens of SAE circuit, we validate conclusions in previous work and also discover some subtle mechanistic distinction in their corresponding circuits.
@@ 5.1 SAE Circuits Closely Agree with Head-Level Ones
We manage to find the end-to-end information flow in the IOI task example $s_{\text{Mary}}$ and its variant $s_{\text{John}}$ with Hierarchical Attribution. Then we identify the pivotal attention heads in the isolated subgraph and attribute their QK scores to earlier SAE features. Discovered SAE feature circuits are of strong consistency with those found on the basis of attention heads: (1) Name Mover features correspond to Name Mover Heads (L9A.H6, L9A.H9); (2) Association features correspond to S-Inhibition Heads (L7A.H3, L7A.H6, L8A.H10); (3) Induction features correspond to Induction Heads (L5A.H5, L6A.H9); (4) Preceding features correspond to Previous Token Heads (L2A.H2, L3A.H2, L4A.H1).
@@ 5.2 Zooming in on SAE Circuits Yields New Discoveries
We present a concrete example in the wild that SAE circuits convey more information than their coarse-grained counterparts. We believe this is a positive signal for us to obtain deeper understanding of language model circuits. Despite the consistency of involved attention heads in $s_{\text{John}}$ and $s_{\text{Mary}}$, these two circuits are actually composed of completely different SAE features, as shown in Figure 6(b).
We start with interpreting how GPT2 predicts " Mary" given the prompt "When John and Mary went to the store, John gave the bag to" ($s_{\text{John}}$). Though greatly simplified, the information flow is still somehow complicated. We further pick four pivotal feature clusters, as marked in Figure 6. A non-rigorous interpretation of them is as follows.
@@
    S2 i.e. the second "John" activates an induction feature. It enhances the logit of "and" though its next token is not.
@@
The interpretation above is highly dependent on the fact that the Indirect Object is present after an " and". However, things are quite different in $s_{\text{Mary}}$ since it comes before the "and". In fact, token " Mary" first activates a Center Entity feature, whose explanation given by GPT-4 is "People or Objects that is likely to be the main topic of the article". The last token still seeks to associate a previously occurring entity but is informed to retrieve the Center Entity instead since the Consecutive Entity Association feature has been suppressed by repeated " John"s.
@@ 6 Related Work
Mechanistic Interpretability (Olah et al., 2020b; a) deems model components e.g. attention heads and MLP neurons as primitives and explain how they interact with model input and output. This line of research has succeeded in identifying attention-based circuits implementing various NLP tasks (Olsson et al., 2022; Wang et al., 2023; Stefan Heimersheim, 2023). Efforts are also made to interpret polysemantic MLP neurons (Gurnee et al., 2023) and editing information stored in MLP parameters (Meng et al., 2022; Sharma et al., 2024).
@@
Previous work mechanistically interprets circuits connecting attention heads and MLP neurons (Olsson et al., 2022; Wang et al., 2023; Conmy et al., 2023). As for SAE circuits, He et al. (2024) makes a linear approximation of MLP layers by fixing the gate mask of the non-linear activation function; Marks et al. (2024) estimate the indirect effect of each SAE feature with attribution patching (Kramár et al., 2024), which also makes linear assumption of non-linear functions. In contrast, we refactor our computation graph to be completely linear w.r.t. OV and MLP circuits without approximation.
@@ 7 Conclusion and Limitation
We frame a pipeline to identify fine-grained circuits in Transformer language models. With Sparse Autoencoders and Skip SAEs, we refactor the model’s computation to linear (with respect to a single input). We also propose an efficient approach to isolate subgraphs (i.e. circuits). We showcase that finer-grained circuit analysis provides more beautiful and detailed structures in Transformers. One limitation of our work is that our analysis is specific to certain inputs and might not generalize to other settings. We deem this as a trade-off between granularity and universality. Some extension can be made to extract more general circuits in regard to more abstract behaviors. We leave this for future work.
@@ Appendix A Sparse Autoencoder Training
We trained an SAE (Section 2.1) on each of the outputs of the 12 attention layers and 24 residual stream activation (before entering attention layers and MLP layers). We trained a Skip SAE (Section 2.2) through each MLP layer, using residual stream activation before MLP as input and MLP output activation as label. Here’s our training settings:
@@
  • We use the Adam optimizer with a learning rate of 4e-4 and betas of (0, 0.9999) for 1 billion tokens from the OpenWebText (Radford et al., 2019) corpus. We trained against a reconstruction loss (measured by MSE of input and reconstructed output), a sparsity loss (proximated by the L1 norm of the feature activations, with a coefficient of 8e-5 (1.2e-4 for attention output SAEs)), and a ghost gradient loss. A batch size of 4,096 is used. We use an NVIDIA A100-80GB GPU for training of each SAE, which lasts for 20 hours.
  • First 256 tokens of each sequence are used as input, discarding the rest tokens and sequences shorter than 256 tokens. Generated activations are shuffled actively in an activation buffer.
  • We normalize the input activations to have a norm of the square root of LM hidden size (i.e. $\sqrt{768}$ for GPT-2 Small). We further normalize the MSE loss by the variance of output along the hidden dimension (a bit like the latter part in LayerNorm, except that we’re not taking the mean of output).
@@
  • We use untied weights for the encoder and decoder. Decoder bias (or pre-encoder bias) is removed (for the sake of simplier circuit analysis). Decoder norm are reset to less than or equal to 1 after each training step.
  • *We prune the dictionary features with norm less than 0.99, max activation less than 1, and activation frequency less than 1e-6 after training.
@@ A.1 Feature Pruning
Some of the SAE features obtained from end-to-end training are too sparse (i.e. can hardly be activated) to reflect a certain aspect of the input corpus. These features are more like "local codes" (in neuroscience). They are activated by very specific tokens. These features are trivial and not helpful for understanding an activation pattern from a compositional perspective. Feature pruning aims to remove these trivial features and keep the more meaningful ones.
@@
In SAE training, we use an L1 loss as a differentiable approximation of L0 loss, to encourage sparsity in the feature activations. The side effect is that the L1 loss as well encourages a lower value of the feature activations, and a larger feature norm. Thus, if a feature is really "useful" in reconstructing the input, it should have a norm as large as possible. We prune the features without the tendency to grow.
@@
Given a fixed norm of the feature, a feature with a low max activation value makes little contribution to reconstructing the input. We find this kind of features activated in some non-related situations, and thus non-interpretable. We empirically set the threshold to 1 and prune the features below it.
@@
A feature with an ultra-low activation frequency is considered to be too local to be useful. We find this kind of features often correspond to some specific tokens in some specific contexts, which is too trivial to be recognized as a feature. We empirically set the threshold to 1e-6 and prunes the features activated in a frequency below it.
@@ A.2 Finetuning against Feature Suppression
To address this issue, we finetune the decoder and a feature activation scaler of the pruned SAEs on the same dataset. Only the reconstruction loss (i.e. the MSE loss) is applied in this finetuning process. Encoder weights are fixed during this process to keep sparsity of the dictionary. Finetuning may also repair flaws introduced in the pruning process, and improve the overall reconstruction quality.
@@ A.3 Statistics of Sparse Autoencoders
We evaluate the L0 loss, variance explained and reconstruction CE loss of each trained SAEs. The L0 loss computes the average feature activated at each token. Variance explained computes
@@
which is a measurement of the proportion to which an SAE accounts for the activation variation. Reconstruction CE loss is the final cross-entropy loss of the language model where the activation is replaced with the SAE reconstructed one. Reconstruction CE score shows how good the reconstruction CE loss is w.r.t the original CE loss and the ablated CE loss, by computing
@@
Table 2: Statistics of MLP Skip SAEs
@@ Appendix B General Direct Contribution Computation
Figure 7: The workflow of interpreting a non-linear transformation where the transformation matrix can be linearly decomposed. We first compute the direct contribution $W^{\prime}v_{i}$ to the transformation matrix $W$ of each partition $v_{i}$ of $x$ so as to reveal the formation of $W$, and then treat the computed $W$ as constant to compute the final direct contribution $Wv_{i}$.
Nevertheless, computation in practical neural networks is often much complicate than the above affine transformation or its simple nesting. Non-linear transformation (e.g. LayerNorm, Softmax, ReLU) is ubiquitous. We cannot simply ignore these non-linear operators since the powerful fitting capacity of neural networks often just come from the non-linear parts. To deal with these non-linear transformations, we propose a more general direct contribution computing strategy. For any transformation $f:\mathbb{R}^{H}\to\mathbb{R}^{K}$ where $f$ has a form of $f(x)=W(x)x+b$, $W:\mathbb{R}^{H}\to\mathbb{R}^{K\times H}$, $b\in\mathbb{R}^{K}$, we have
@@
where we treat $W(x)$ as a constant linear transformation matrix. Then we can claim that $i$-th partition $v_{i}$ contributes to the result $f(x)$ by $W(x)v_{i}$ through the posterior linear transformation with a constant $W(x)$. We must state this contribution computation is nothing but trivial if we don’t further interpret how partitions affect $W(x)$ and the relating impact to the following transformation, or further restrict the $W(x)$to make sure it’s just close to a constant or its variation is unimportant. Thus, for $W$ that having a similar form as $f$, e.g. $W(x)=W^{\prime}(x)x+B$, $W^{\prime}:\mathbb{R}^{H}\to\mathbb{R}^{(K\times H)\times H}$, $B\in\mathbb{R}^{K\times{}H}$, we can iteratively apply the linear decomposition Eq. 4 to $W$ (which we use to interpret attention pattern in Sec. 2.3),
@@
As a brief summary, the core idea of direct contribution computation for any non-linear function is to first compute how the non-linear part is formed w.r.t each input partition by iteratively applying direct contribution computation, and then consider the non-linear part as determined, regard the function to be linear, and compute a linear contribution to the function output. We usually allow the determined non-linear part to go through a simple extra activation function like Softmax or ReLU, since this will not undermine the understanding of this non-linear part. This workflow can be applied to non-linear function like bi-linear function and attention.

### v2

@@ 4 Attributing Intermediate SAE Features
An exciting application of Sparse Autoencoders is that they serve as unsupervised feature extractors in the vast hidden activation space. This opens up opportunities for understanding intermediate activations and local circuit discovery, i.e., identifying a subgraph activating a given SAE feature instead of end-to-end circuits.
@@ 4.1 How Transformers Implement In-Bracket Features
We start from a series of In-Bracket features in attention blocks of early layers, which activate on tokens inside of brackets, e.g., deactivated [activated] deactivated. These features will demonstrate higher activation in deeper nesting of brackets, imitating the behavior of finite state automata (Bricken et al., 2023) with states of bracket nesting hierarchy. We find an In-Square-Bracket feature and an In-Round-Bracket feature in SAEs trained on layer 1 attention block output, which we call L1A throughout this paper. Since they are at rather early layers, we leverage our Direct Contribution analysis to see how earlier features produce them.
Open-bracket features activate in-bracket ones. Figure 3(a) illustrates a simple two-layer bracket circuit in the wild. We inspect contributions to the In-Square-Bracket feature in a template, e.g. "0 0 [1 1 1 [2] 3] 4", at token "1"s, "2", " 3" and " 4". Experiments show that the activation is mainly promoted by an L0M feature activated by the token " [". It takes on 104.1%, 102.6% and 314.2% of the In-Square-Bracket feature’s activation respectively at token "1", "2", and " 3", respectively. An average of 83.8% of these contributions comes through the attention head 1 of L1A, i.e., L1A.H1.
@@ 4.2 Revisiting Induction Behavior from the SAE Lens
Induction Heads (Olsson et al., 2022) is an important type of compositional circuit with two attention layers which try to repeat any 2-gram that occurred before, i.e. [A][B] … [A] -> [B]. These circuits are believed to account for most in-context learning functionality in large transformers. Compared to the massive existing literature in understanding the induction mechanism in the granularity of attention heads (Olsson et al., 2022; Hendel et al., 2023; Ren et al., 2024), inter alia, we seek to present a finer-grained level interpretation of such behavior.
Induction features form a huge feature family. These features are found to be identified by the logit of tokens they enhance through the logit lens (nostalgebraist, 2020). We first study a Capital Induction feature contributing to logits of single capital letters on a curated input "Video in WebM support: Your browser doesn’t support HTML5 video in WebM." (Figure 4(a)). This feature is activated on the second " Web" and amplifies the prediction of "M", copying its previous occurrence.
@@
Figure 4: "Web"(L0M.1270 and L1M.23399) and "Web" Preceding features (L2A.14876 and L2A.17608) jointly lead to QK attention of an induction head. The "M" feature is copied to the last token for the next token prediction.
Upstream Contribution through OV Circuit We notice that a series of "M" features in the residual stream of the first "M" constitute most of the Capital Induction feature’s activation through OV circuits. L0M.88 takes the lead, which contributes 35.0% of the feature activation. Auxiliary features from L0A, L1M, and L3M either directly indicate the current token as "M" or indicate the current token as a single capital letter. Top 7 of the auxiliary features account for another 33.0% of the feature activation. Most of these contributions come from L5A.H1, which we along with a concurrent research (Krzyzanowski2024attention_saes) identify as an induction head.
Upstream Contribution to QK Attention To study how this induction head attends to the first "M", we attribute the attention score to upstream feature pairs. The commonality of top contributors is a " Web" feature attending to a " Web" Preceding feature (i.e., its previous token is " Web"), as shown in Figure 4(b).
Attributing Preceding features We further study how " Web" Preceding features indicate previous tokens. These contributions mainly come through L2A.H2, which we think to be a previous token head. The relatively high attention score for the previous token can be attributed to a group of L0A features collecting information from positional embeddings.
@@ 5 Revisiting Indirect Object Identification Circuits from the SAE Lens
Figure 5: In $s_{\text{John}}$, the consecutive entity feature (denoted as A in Figure 5(a)) serves as the key vector for Name Mover Heads to attend to and copy the answer entity to the last token’s residual stream. Such a mechanism does not work in $s_{\text{Mary}}$ because the correct answer is no longer a consecutive entity (i.e., the entity present after the token and). See Appendix E for a detailed interpretation of these two examples.
For end-to-end circuits in GPT-2 Small, we choose to investigate a task called Indirect Object Identification (IOI) (Wang et al., 2023) with Hierarchical Attribution. For instance, GPT-2 can predict " Mary" following the prompt "When Mary and John went to the store, John gave the bag to". We call this prompt $s_{\text{Mary}}$ since it starts with " Mary" and a variant $s_{\text{John}}$ with a swap in the first two names, i.e. "When John and Mary went to the store, John gave the bag to". The answer to both prompts is " Mary", which GPT-2 is able to predict. Existing literature studying this problem does not distinguish between these two templates. Through the lens of SAE circuits, we validate conclusions in previous work and also discover some subtle mechanistic distinctions in their corresponding circuits.
@@ 5.1 SAE Circuits Closely Agree with Head-Level Ones
We manage to find the end-to-end information flow in the IOI task example $s_{\text{Mary}}$ and its variant $s_{\text{John}}$ with Hierarchical Attribution. Then, we identify the pivotal attention heads in the isolated subgraph and attribute their QK scores to earlier SAE features. Discovered SAE feature circuits are of strong consistency with those found based on attention heads: (1) Name Mover features correspond to Name Mover Heads (L9A.H6, L9A.H9); (2) Association features correspond to S-Inhibition Heads (L7A.H3, L7A.H6, L8A.H10); (3) Induction features correspond to Induction Heads (L5A.H5, L6A.H9); (4) Preceding features correspond to Previous Token Heads (L2A.H2, L3A.H2, L4A.H1).
@@ 5.2 Zooming in on SAE Circuits Yields New Discoveries
We present a concrete example in the wild that SAE circuits convey more information than their coarse-grained counterparts. We believe this is a positive signal for us to obtain a deeper understanding of language model circuits. Despite the consistency of involved attention heads in $s_{\text{John}}$ and $s_{\text{Mary}}$, these two circuits are actually composed of completely different SAE features, as shown in Figure 5(b).
We start with interpreting how GPT-2 predicts " Mary" given the prompt "When John and Mary went to the store, John gave the bag to" ($s_{\text{John}}$). Though greatly simplified, the information flow is still somehow complicated. We further pick four pivotal feature clusters, as marked in Figure 5. A non-rigorous interpretation of them is as follows.
@@
    S2, i.e., the second "John" activates an induction feature. It enhances the logit of "and" though its next token is not.
@@
The interpretation above highly depends on the fact that the Indirect Object is present after an " and". However, things are quite different in $s_{\text{Mary}}$ since it comes before the "and". In fact, token " Mary" first activates a Center Entity feature, whose explanation given by GPT-4 is "People or Objects that is likely to be the main topic of the article". The last token still seeks to associate a previously occurring entity but is informed to retrieve the Center Entity instead since the Consecutive Entity Association feature has been suppressed by repeated " John"s.
@@ 6 Related Work
Mechanistic Interpretability (Olah et al., 2020b; a) deems model components, e.g., attention heads and MLP neurons, as primitives and explains how they interact with model input and output. This line of research has succeeded in identifying attention-based circuits implementing various NLP tasks (Olsson et al., 2022; Wang et al., 2023; Stefan Heimersheim, 2023). Efforts are also made to interpret polysemantic MLP neurons (Gurnee et al., 2023) and editing information stored in MLP parameters (Meng et al., 2022; Sharma et al., 2024).
@@
Previous work mechanistically interprets circuits connecting attention heads and MLP neurons (Olsson et al., 2022; Wang et al., 2023; Conmy et al., 2023). As for SAE circuits, He et al. (2024) makes a linear approximation of MLP layers by fixing the gate mask of the non-linear activation function; Marks et al. (2024) estimates the indirect effect of each SAE feature with attribution patching (Kramár et al., 2024), which also makes linear assumption of non-linear functions. In contrast, we refactor our computation graph to be completely linear w.r.t. OV and MLP circuits without approximation.
@@ 7 Conclusion and Limitation
We frame a pipeline to identify fine-grained circuits in Transformer language models. With Sparse Autoencoders and Transcoders, we refactor the model’s computation to linear (with respect to a single input). We also propose an efficient approach to isolate subgraphs (i.e. circuits). We showcase that finer-grained circuit analysis provides more beautiful and detailed structures in Transformers. One limitation of our work is that our analysis is specific to certain inputs and might not generalize to other settings. We deem this as a trade-off between granularity and universality. Some extensions can be made to extract more general circuits regarding more abstract behaviors. We leave this for future work.
@@ Appendix A Sparse Autoencoder Training
We trained an SAE (Section 2.1) on each of the outputs of the 12 attention layers and 24 residual stream activation (before entering attention layers and MLP layers). We trained a Skip SAE (Section 2.2) through each MLP layer, using residual stream activation before MLP as input and MLP output activation as the label. Here are our training settings:
@@
  • We use the Adam optimizer with a learning rate of 4e-4 and betas of (0, 0.9999) for 1 billion tokens from the OpenWebText corpus. We trained against a reconstruction loss (measured by MSE of input and reconstructed output), a sparsity loss (proxied by the L1 norm of the feature activations, with a coefficient of 8e-5 (1.2e-4 for attention output SAEs)), and a ghost gradient loss. A batch size of 4,096 is used. We use an NVIDIA A100-80GB GPU for training of each SAE, which lasts for 20 hours.
  • The first 256 tokens of each sequence are used as input, discarding the remaining tokens and sequences shorter than 256 tokens. Generated activations are shuffled actively in an activation buffer.
  • We normalize the input activations to have a norm of the square root of LM hidden size (i.e., $\sqrt{768}$ for GPT-2 Small). We further normalize the MSE loss by the variance of output along the hidden dimension (a bit like the latter part in LayerNorm, except that we’re not taking the mean of output).
@@
  • We use untied weights for the encoder and decoder. Decoder bias (or pre-encoder bias) is removed (for the sake of simpler circuit analysis). Decoder norms are reset to less than or equal to 1 after each training step.
  • *We prune the dictionary features with a norm less than 0.99, max activation less than 1, and activation frequency less than 1e-6 after training.
@@ A.1 Feature Pruning
Some of the SAE features obtained from end-to-end training are too sparse (i.e., can hardly be activated) to reflect a certain aspect of the input corpus. These features are more like "local codes" (in neuroscience). They are activated by very specific tokens. These features are trivial and not helpful for understanding an activation pattern from a compositional perspective. Feature pruning aims to remove these trivial features and keep the more meaningful ones.
@@
In SAE training, we use an L1 loss as a differentiable approximation of L0 loss, to encourage sparsity in the feature activations. The side effect is that the L1 loss as well encourages a lower value of the feature activations and a larger feature norm. Thus, if a feature is really "useful" in reconstructing the input, it should have a norm as large as possible. We prune the features without the tendency to grow.
@@
Given a fixed norm of the feature, a feature with a low max activation value contributes little to reconstructing the input. We find this kind of feature activated in some non-related situations and thus non-interpretable. We empirically set the threshold to 1 and prune the features below it.
@@
A feature with an ultra-low activation frequency is considered too local to be useful. We find that these features often correspond to some specific tokens in some specific contexts, which is too trivial to be recognized as a feature. We empirically set the threshold to 1e-6 and prune the features activated at a frequency below it.
@@ A.2 Finetuning against Feature Suppression
To address this issue, we finetune the decoder and a feature activation scaler of the pruned SAEs on the same dataset. Only the reconstruction loss (i.e., the MSE loss) is applied in this fine-tuning process. Encoder weights are fixed during this process to keep the sparsity of the dictionary. Finetuning may also repair flaws introduced in the pruning process and improve the overall reconstruction quality.
@@ A.3 Statistics of Sparse Autoencoders
We evaluate the L0 loss, variance explained, and reconstruction CE loss of each trained SAE. The L0 loss computes the average feature activated at each token. Variance explained computes
@@
which measures the proportion to which an SAE accounts for the activation variation. Reconstruction CE loss is the final cross-entropy loss of the language model, where the activation is replaced with the SAE reconstructed one. The reconstruction CE score shows how good the reconstruction CE loss is w.r.t the original CE loss and the ablated CE loss by computing
@@
Table 2: Statistics of MLP Transcoders
@@ Appendix B General Direct Contribution Computation
Figure 6: The workflow of interpreting a non-linear transformation where the transformation matrix can be linearly decomposed. We first compute the direct contribution $W^{\prime}v_{i}$ to the transformation matrix $W$ of each partition $v_{i}$ of $x$ to reveal the formation of $W$, and then treat the computed $W$ as constant to compute the final direct contribution $Wv_{i}$.
Nevertheless, computation in practical neural networks is often much complicated than the above affine transformation or its simple nesting. Non-linear transformation (e.g. LayerNorm, Softmax, ReLU) is ubiquitous. We cannot simply ignore these non-linear operators since the powerful fitting capacity of neural networks often just comes from the non-linear parts. To deal with these non-linear transformations, we propose a more general direct contribution computing strategy. For any transformation $f:\mathbb{R}^{H}\to\mathbb{R}^{K}$ where $f$ has a form of $f(x)=W(x)x+b$, $W:\mathbb{R}^{H}\to\mathbb{R}^{K\times H}$, $b\in\mathbb{R}^{K}$, we have
@@
where we treat $W(x)$ as a constant linear transformation matrix. Then, we can claim that $i$-th partition $v_{i}$ contributes to the result $f(x)$ by $W(x)v_{i}$ through the posterior linear transformation with a constant $W(x)$. We must state this contribution computation is nothing but trivial if we don’t further interpret how partitions affect $W(x)$ and the related impact to the following transformation or further restrict the $W(x)$to make sure it’s just close to a constant or its variation is unimportant. Thus, for $W$ that having a similar form as $f$, e.g. $W(x)=W^{\prime}(x)x+B$, $W^{\prime}:\mathbb{R}^{H}\to\mathbb{R}^{(K\times H)\times H}$, $B\in\mathbb{R}^{K\times{}H}$, we can iteratively apply the linear decomposition Eq. 7 to $W$ (which we use to interpret attention pattern in Sec. 2.3),
@@
As a brief summary, the core idea of direct contribution computation for any non-linear function is to first compute how the non-linear part is formed w.r.t each input partition by iteratively applying direct contribution computation, and then consider the non-linear part as determined, regard the function to be linear, and compute a linear contribution to the function output. We usually allow the determined non-linear part to go through a simple extra activation function like Softmax or ReLU, since this will not undermine the understanding of this non-linear part. This workflow can be applied to non-linear functions like bi-linear functions and attention.